Past Meetings


state college snowfall state college snowfall

Unviersity Park, PA

May 25-26, 2017


HET Board Meeting Agenda (pdf)

December Minutes (pdf) [12/02/16]

Science Operation Presentation on Dec - March & Current Trimester - Shetrone (pdf) [05/22/17]

Science Operations Report on Dec - March & Current Trimester - Shetrone (pdf) [05/22/17]

HET Facility Managers Report - Kriel (pdf) [05/22/17]

Major Issues of Facility Manager's Report - Kriel (pdf) [05/22/17]

Current Telescope Status - Kriel (pdf) [05/22/17]

HET SCS Upgrade and Priority - Kriel (pdf) [05/22/17]

Major Issues & Discussion of Telescope Systems Performance Report - Drory (pdf) [06/07/17]

Major Issues & Discussion LRS2 Performance Update Report - Hill (pdf) [05/30/17]

VIRUS: Current Status - Hill (pdf) [05/30/17]

Major Steps to Implement HRS Operations - MacQueen (pdf) [05/30/17]

HPF Status, Schedule & Plan for Commissioning Report - Mahadevan (pdf) [05/30/17]

HET Scientist & User Committee Report - Schneider (pdf) [05/22/17]

Parallel Science Updates - Shetrone (pdf) [05/22/17]

HET Data Management & HETDEX Commissioning Progress - Gabhardt (pdf) [05/22/17]

HET Data Scientist Status - Armandroff (pdf) [05/22/17]

Operating Budgets -Status FY 16-17 & Proposed FY 17-18 - Tarver (pdf) [05/30/17]

LRS2 Science Updates - Hill (pdf) [05/30/17]

HETDEX Science Updates - Ciardullo (pdf) [05/30/17]

This is an abbreviation version of the Past Meetings archive. To access full archived version, navigate to the full archive version. (Login required.)